We all believe that we see most of what goes on around us, at least the most obvious things Recently, however, psychologists have identified a phenomenon called "selective looking" which means that, if our attention is drawn to one aspect of what we see, we can miss even seemingly obvious features presented at the same time In a striking demonstration of this phenomenon, a seres of randomly chosen students was shown a video of several people throwing a basketball around, and they were asked to count how many times the people in white shirts threw the ball(Simons and Chabris 1999). In the middle of the video, a woman dressed as a gorilla walked through the shot, pausing in the center to thump her chest, and then walked out of the shot. Of the 12 students watching the video, only five noticed the gorilla.

a. What is the best estimate from these data of the proportion of students in the population who notice the woman in the gorilla suit?

b. What is the 95% confidence interval for the proportion of students in the population who notice the woman in the gorilla suit. What are the lower and upper bounds? (Sample Mean - 2*Standard Error of the Sample Mean)
Answer #1

a)
pcap = 5/12 = 0.4167

b)
SE = sqrt(0.4167*(1-0.4167)/12) = 0.1423

ME = 1.96*0.1423 = 0.2789

CI = (0.4167 - 0.2789, 0.4167 + 0.2789)
= (0.1378, 0.6956)